# Recoverable strain from reverse plasticity

### v1

@@ Abstract
Recoverable strain is the strain recovered once a stress is removed from a body, in the direction opposite to that in which the stress had acted. To date, the phenomenon has been understood as being elastic in origin: polymer chains stretched in the direction of an imposed stress will recoil after the stress is removed, for example. Any unrecoverable strain is instead attributed to irreversible plastic deformations. Here we study theoretically recoverable and unrecoverable strain within the soft glassy rheology model, aimed at describing the rheology of elastoplastic yield stress fluids and amorphous soft solids. We consider a material subject to the switch-on of a shear stress that is held constant before later being set back to zero, after which the strain recovery is observed. After an initially fast recoil that is indeed elastic in nature, consistent with the usual intuition, we find that significant subsequent strain recovery then arises not via further recoil of elasticity but instead by ‘reverse plasticity’, in a sense that we discuss carefully. Also unexpectedly, although in rare parameter regimes, this plastic part of the strain post switch-off does not always in fact recover in the negative direction, counter to that of the previously imposed stress, but can sometimes continue to accumulate in the forward direction. The recovery is then non-monotonic overall, reminiscent of recent observations of non-monotonic stress relaxation after straining, which have been attributed to complex material memory effects. We elucidate the mechanisms underlying these behaviours in terms of the evolution of the SGR model’s population of elastoplastic elements, and suggest connections with the notion of ‘reversible plasticity’, as discussed in the recent physics literature.
@@ I Introduction
Counterintuitively, we find that a significant component of strain recovery post switch-off arises not via the recoil of elasticity, but instead by reverse plasticity, in a sense that we discuss carefully. This observation holds robustly across broad parameter regimes. Unexpectedly – although in relatively rare parameter regimes – we also find that the plastic strain post switch-off does not always recover in the negative direction, counter to that of the originally imposed stress, but can sometimes continue to accumulate in the forward direction. The recovery process is then non-monotonic overall, with the initially backward elastic recoil followed by a (lesser) forward plastic straining. This is reminiscent of recent observations of non-monotonic stress relaxation after straining Hendricks et al. (2019); Sudreau et al. (2022); Murphy et al. (2020); Mandal et al. (2021), which have been attributed to complex material memory effects Keim et al. (2019). We provide a mechanistic understanding of these observations in terms of the evolution of the SGR model’s population of elastoplastic elements. We also discuss possible connections with the concept of reversible plasticity, which has gained increasing attention in the recent physics literature.
Indeed, a key concept in the rheology of amorphous materials is that of intermittent local plastic events, sometimes called T1 events, arising at soft spots or shear transformation zones Falk and Langer (2011), coupled by elastic stress propagation Picard et al. (2004). Within this framework, oscillatory shear experiments on jammed particle rafts identified three different regimes of imposed strain amplitude Galloway et al. (2020); Keim and Arratia (2014, 2013). At low amplitudes, the material responds elastically, with no plastic events. At high amplitudes, irreversible plastic events occur within each cycle, with significant cycle-to-cycle stroboscopic changes in particle positions. More surprisingly, at intermediate strain amplitudes these experiments report reversible local plastic events within each cycle. Such events dissipate energy, consistent with the notion of plasticity, but exactly reverse in the opposite part of the cycle, with zero stroboscopic cycle-to-cycle change in particle positions. Reversible plastic events have also been reported in experiments on foams Lundberg et al. (2008), in particle simulations Lundberg et al. (2008); Regev et al. (2013); Priezjev (2016); Szulc et al. (2022); Nagasawa et al. (2019) and in elastoplastic models Khirallah et al. (2021). Reversible plastic limit cycles have likewise been studied in models of hysterons Keim and Paulsen (2021) and iterated maps Mungan and Witten (2019). For a review, see Ref. Reichhardt et al. (2022).
@@
The paper is structured as follows. In Sec. II we outline the SGR model Sollich et al. (1997); Sollich (1998), within which we shall study the phenomenon of strain recovery. The step stress protocol that we shall simulate is specified in Sec. III. In Sec. IV we define our units, summarise the parameters of the model and protocol, and discuss the rheological quantities to be measured. Our results are presented in Sec. V. We set out our conclusions in Sec. VI.
@@ II Soft glassy rheology model
The soft glassy rheology (SGR) model Sollich et al. (1997); Sollich (1998) considers the dynamics of an ensemble of elastoplastic elements that explore a landscape of energy traps. Each element is notionally taken to represent a local mesoscopic region of a soft glassy material: a few tens of droplets in a dense emulsion, for example. Each element is assumed small relative to any macroscopic variations in the flow field, on the one hand, yet large enough to allow the definition of local continuum variables of shear strain $l$ and shear stress $kl$, on the other. The local strain variable $l$ is taken to describe the element’s state of elastic deformation relative to a locally undeformed equilibrium. Between local plastic yielding events, defined below, the strain of each element is assumed elastically to follow the macroscopically imposed strain rate $\dot{\gamma}$, such that $\dot{l}=\dot{\gamma}$.
Plasticity is incorporated by assuming that these local stresses are intermittently released by local plastic yielding events. In each such yielding event, a mesoscopic region of material is assumed suddenly to rearrange itself into a new local configuration, relative to a new state of local equilibrium. This is modelled by the corresponding representative elastoplastic element hopping between traps in the model’s energy landscape: at any instant in time, an element that lies in a trap of energy depth $E$ and has local shear strain $l$ is assumed to have a probability per unit time of hopping given by
$r(E,l)=\tau_{0}^{-1}{\rm max}\left\{1,\exp\left[(\tfrac{1}{2}kl^{2}-E)/x\right]\right\},$ (1)
@@
in which $x_{\rm g}$ is the model’s glass transition (effective) temperature, discussed further below. The element is further assumed to reset its local strain $l$ to zero, so relaxing its stress. (Post-hop frustration can by incorporated by assuming a post-hop strain distribution of non-zero width, though we do not consider that here.)
@@
The advection term on the left hand side captures the elastic loading of each element within its trap. The first term on the right hand side describes the local plastic relaxation events described above. Because these are modelled by hops out of traps, this takes the form of a ‘death’ term. The second term on the right hand side is a ‘rebirth’ term, describing the hopping of freshly yielded elements into the bottom of new traps, $l=0$, with the new trap depth selected randomly from the prior $\rho(E)$ of Eqn. 2. The ensemble average hopping rate
@@ III Protocol
Sample preparation prior to shear is modelled by assuming a sudden deep quench at some time $t=-t_{\rm w}$, from infinite effective temperature to the final working value $x$, which is held constant thereafter. For times $-t_{\rm w}<t<0$ the sample is allowed to age undisturbed, without any strain or stress being applied. Throughout most of the paper, we assume that all elemental strains are equal to zero during this initial ageing stage, $l=0$. An exception can be found in Figs. 3 and 4, where we assume an initial Gaussian distribution of $l$ values, of narrow standard deviation $l_{0}=0.05$. This does not affect the basic physics of interest in this work, but merely allows us to show as a narrow Gaussian in Fig. 4 what would otherwise be an unplottable initial delta function distribution.
@@ IV Units, parameter values and measured quantities
To solve the SGR model numerically, we simulate a population of $M$ SGR elements. In any timestep $dt$, any element is hopped with probability $rdt$, with $r$ that element’s probability per unit time of hopping as given by Eqn. 1 above.
@@
At the instant of stress switch-off, all elements instantaneously strain elastically backwards, giving a global strain change $\Delta\gamma=-\sigma_{0}$ (in our units). We then measure the subsequent strain evolution as a function of the time $t-t_{\rm stop}^{+}$ since switch-off. Our interest in particular will be in the total reverse strain $\Delta\gamma_{\rm rec}$ accumulated in the long time limit, $t-t_{\rm stop}\to\infty$, beyond that which arose elastically during the reverse step. As noted above, in rare cases we shall find $\Delta\gamma_{\rm rec}<0$: the plastic straining after the initial reverse elastic recoil in fact takes place in the forward direction.
@@ V.1 Forward strain during stress application
Following the application of a step stress of magnitude $\sigma_{0}$ at time $t=0$, a material’s creep response is characterised by the strain curve $\gamma(t)$. As noted above, in the SGR model this shows an initially elastic step strain of magnitude $\gamma_{0}=\sigma_{0}/k\;(=\sigma_{0}$ in our units) at $t=0$, followed by subsequent plastic straining for $t>0$. Fig. 1a) shows this plastic part of the creep, beyond the initial elastic step, normalised by the stress amplitude $\sigma_{0}$. Results are shown for a fixed sample age $t_{\rm w}$, and several imposed stresses in curves upwards. The corresponding strain rate curves $\dot{\gamma}(t)$ are shown in panel b).
For the noise temperature explored here, $x=0.3$, deep within the glass phase, the SGR model’s yield stress $\sigma_{\rm Y}\approx 0.748$. For imposed stresses $\sigma_{0}<\sigma_{\rm Y}$, the strain rate decreases perpetually as a power law $\dot{\gamma}\sim t^{-\alpha}$. The exponent $10<\alpha<$ 1 depends on the noise temperature $x$. The strain correspondingly increases sublinearly as $\gamma\sim t^{1-\alpha}$, corresponding to sustained power-law creep. In contrast, for imposed stresses $\sigma_{0}>\sigma_{\rm Y}$, an initial interval of creep (for $\sigma_{0}\gtrapprox\sigma_{\rm Y}$ at least) is interrupted by yielding, in which the strain rate curves upwards before finally settling to a time-independent value prescribed by the steady state flow curve $\sigma_{\rm ss}(\dot{\gamma})$, with $\sigma_{\rm ss}-\sigma_{\rm Y}(x)\sim\dot{\gamma}^{1-x}$. The time delay before yielding scales as the sample age $t_{\rm w}$, with a prefactor that diverges as $\sigma_{0}-\sigma_{\rm Y}\to 0^{+}$ Fielding et al. (2000).
@@
Fig. 2b) shows the same data as in Fig. 1, but with the scaled strain rate $\dot{\gamma}(t)/\sigma_{0}$ now plotted parameterically as a function of the forward plastic strain $(\gamma(t)-\gamma_{0})/\sigma_{0}$. Accordingly, time $t$ is not explicitly represented in Fig. 2. The scaled strain on the abscissa however increases monotonically with increasing $t$. The counterpart data for sample ages $t_{\rm w}=10^{1}$ and $10^{6}$ are shown in Figs. 2a) and c) respectively. As can be seen, for imposed stresses above the yield stress, $\sigma_{0}>\sigma_{\rm Y}$, yielding typically occurs at a scaled strain $(\gamma-\gamma_{0})/\sigma_{0}=O(1)$, although the value increases slowly with $t_{\rm w}$ and diverges as $\sigma_{0}\to\sigma_{\rm Y}^{+}$.
A dot at each of several logarithmically spaced values of the abscissa in Fig. 2 shows the values of scaled forward plastic strain $(\gamma-\gamma_{0})/\sigma_{0}=\Delta\gamma_{\rm f}/\sigma_{0}$ at which we shall later, in Fig. 5, explore the amount of strain $\Delta\gamma_{\rm rec}$ that is ultimately recovered after the stress is switched off.
@@ V.2 Strain recovery after stress switch-off
Having mapped out the forward creep and/or yielding and flow that arise while the stress is held imposed, we now explore the strain recovery after the stress is switched off. Our focus will be in particular on the part of this recovery that stems from reverse plasticity, after the initial elastic recoil at the instant of switch-off.
@@
The basic physics that we aim to elucidate is shown in Figs. 3 and 4. The imposed stress $\sigma$ as a function of time $t$ is shown in the top panel of Fig. 3. The corresponding strain response $\gamma(t)$ is shown in the bottom panel. As described above, a stress of amplitude $\sigma_{0}$ is switched on at time $t=0$. (The preparation phase during which the sample is allowed to age undisturbed for a time $t_{\rm w}$ prior to $t=0$ is not shown.) At the instant of switch-on, the system shows an instantaneous forward elastic step strain of amplitude $\gamma_{0}=\sigma_{0}$, in our units in which $k=1$. (In experimental practice, this forward strain would be rapid but not instantaneous, with a rate limited by inertia or dissipation.) The stress is then held constant for some time interval, over which the strain increases further due to plastic yielding events, this forward straining having been explored in Sec. V.1. Once it has thus increased by an amount $\Delta\gamma_{\rm f}$ beyond $\gamma_{0}$, the stress is switched off. This defines the switch-off time $t_{\rm stop}$. At the instant of switch-off, $t=t_{\rm stop}$, the system shows an instantaneous backward elastic step strain, $\Delta\gamma=-\sigma_{0}=-\gamma_{0}$. The stress is then held at zero for all subsequent times $t>t_{\rm stop}$. As a function of the time interval $t-t_{\rm stop}^{+}$, the strain further recovers in the negative direction by an amplitude $\Delta\gamma_{\rm rec}$, beyond the negative elastic step $\Delta\gamma=-\gamma_{0}$ that occurred at the instant of switch-off. This additional recovery $\Delta\gamma_{\rm rec}$, which has a significant amplitude compared with both $\Delta\gamma_{\rm f}$ and $\gamma_{0}$, arises from reverse plasticity post switch-off in a way that we shall now describe.
To understand the physics underlying this strain evolution, let us consider the master equation governing the time evolution of the joint probability distribution of elemental trap depths $E$ and strains $l$. This was specified in Eqn. 3, which we repeat here for convenience:
@@
In contrast, during any interval in which the stress is constant, $\dot{\sigma}=0$, either while the stress of amplitude $\sigma=\sigma_{0}$ is held imposed, or after switch-off when $\sigma=0$, the strain rate is given by
@@
Importantly, this tells us that both the forward straining $\Delta\gamma_{\rm f}$ that occurs after switch-on but before switch-off, and the reverse strain recovery $\Delta\gamma_{\rm rec}$ that occurs after switch-off, are determined purely by the relaxation of local stresses due to local plastic yielding events. Indeed, as these local plastic events occur and elements hop between traps, a global straining must arise in order to advect elements within their traps and exactly counterbalance any such plastic stress loss (or gain), keeping the overall stress constant.
In order fully to understand the forward plastic straining during stress imposition and the reverse plastic strain recovery afterwards, therefore, we must consider in more detail the evolution of the probability distribution of local strains, and the way this determines the strain rate $\dot{\gamma}(t)$ via Eqn. 9. Accordingly, we plot this distribution $P(l,t)$ in Fig. 4 (having first integrated out the trap energy depth variable $E$). We do so at several times $t$ while the stress is held applied (top panel in Fig. 4), and at several times after the stress is switched off (bottom panel). The time at which any distribution is plotted in Fig. 4 is indicated by a cross of corresponding colour in Fig. 3.
@@
Were the strain held constant, with zero strain rate $\dot{\gamma}=0$, the macroscopic consequence of these plastic relaxation events would be a relaxation of the ensemble average stress at rate $\dot{\sigma}=\langle l\,r(E,l)\rangle_{P(t)}$. Conversely, to maintain the stress constant, $\sigma=+\sigma_{0}$ with $\dot{\sigma}=0$, requires a forward strain of corresponding rate $\dot{\gamma}=\langle l\,r(E,l)\rangle_{P(t)}$, ensuring sufficient forward advection of elements within their traps exactly to counterbalance this plastic relaxation. This is the origin of the forward plastic straining seen for times $0<t<t_{\rm stop}$ in Fig. 3. Its consequence is to advect forward at rate $\dot{l}=\dot{\gamma}$ both the spike in probability centred (originally) at $l=\sigma_{0}=\gamma_{0}$ (as it simultaneously depletes due to the plastic relaxation events just described), as well as the newly developing lobe of probability near the origin $l\approx 0$. In this way, the distribution evolves over time from the black to cyan curve in the top panel of Fig. 4.
At the instant the stress is switched off, the distribution of local strains instantaneously shifts by $l\to l-\gamma_{0}$, with $\gamma_{0}=\sigma_{0}$, without changing shape, as each element elastically reduces its strain within its trap. Accordingly, the rightmost (cyan) curve in the top panel of Fig. 4, which obtained at $t=t_{\rm stop}^{-}$, shifts to become the rightmost (magenta) curve in the bottom panel at $t=t_{\rm stop}^{+}$.

### v2

# Recoverable strain in amorphous materials: the role of ongoing plastic events following initial elastic recoil

@@ Abstract
Recoverable strain is the strain recovered once a stress is removed from a body, in the direction opposite to that in which the stress had acted. To date, the phenomenon has been understood as being elastic in origin: polymer chains stretched in the direction of an imposed stress will recoil after the stress is removed, for example. Any unrecoverable strain is instead attributed to irreversible plastic deformations. Here we study theoretically strain recovery within the soft glassy rheology model, aimed at describing the rheology of elastoplastic yield stress fluids and amorphous soft solids. We consider a material subject to the switch-on of a shear stress that is held constant before later being set back to zero, after which the strain recovery is observed. After an initially fast recoil that is indeed elastic in nature, significant further strain recovery then occurs more slowly via the plastic yielding of elements with negative local stresses, opposite to that of the original shear. We elucidate the mechanism that underlies this behaviour, in terms of the evolution of the SGR model’s population of elastoplastic elements. In particular, we show that the initial fast elastic recoil brings to a state of negative local stress those elements that had yielded during the forward straining while the load was applied. The subsequent delayed plastic yielding of these elements with negative stress is the origin of the slow ongoing strain recovery. In this way, counterintuitively, elements that had yielded plastically while the load was applied still contribute significantly to strain recovery after the sample is unloaded. This finding has important consequences for constitutive modeling, because such behaviour can only arise in a constitutive model that evolves a full distribution of local stresses (or multiple moments of such a distribution), rather than a single average stress. Unexpectedly, although in rare parameter regimes, this slow ongoing strain recovery post switch-off does not always in fact recover in the negative direction, counter to that of the previously imposed stress, but can sometimes continue to accumulate in the forward direction. The recovery is then non-monotonic overall, reminiscent of observations of non-monotonic stress relaxation after straining.
@@ I Introduction
After an initially fast recoil that is elastic in nature, we find, counterintuitively, that significant further slow strain recovery then occurs via ongoing plastic events in the reverse direction, opposite to that of the original shear. We provide a mechanistic understanding of this phenomenon in terms of the evolution of the SGR model’s population of elastoplastic elements. In particular, we show that the initial recoil, which elastically advects each element backwards in its trap, leaves in a state of negative stress those elements that had yielded plastically while the material was under load. The subsequent delayed plastic yielding of these elements with negative stress then leads to slow ongoing strain recovery. Because this yielding is of negative stresses, counter to the original shear, we call these “reverse plastic events”. Counterintuitively, therefore, those elements that had yielded plastically while the forward load was applied still make an important contribution to the material’s overall recovery after unloading. This has important consequences for constitutive modeling, because such behaviour can only arise in a constitutive model that evolves a full distribution of local stresses (or multiple moments of such a distribution). A model that instead evolves only a single average stress will be unable to capture it. To illustrate this, we also simulate a simple fluidity model, which shares many features of SGR’s behaviour, including a yield stress, rheological aging, and stress overshoot in shear startup. In evolving only an average stress rather than a stress distribution, however, it lacks the slow ongoing strain recovery predicted by the SGR model.
@@
We show that the observation of slow ongoing strain recovery from reverse plastic events holds robustly across broad parameter regimes. Unexpectedly – although in relatively rare parameter regimes – we also find that the slow ongoing strain post switch-off does not always take place in the negative direction, counter to that of the originally imposed stress, but can sometimes continue to accumulate in the forward direction. The recovery process is then non-monotonic overall, with the initially backward elastic recoil followed by a (lesser) slow forward straining. This is reminiscent of recent observations of non-monotonic stress relaxation after straining Hendricks et al. (2019); Sudreau et al. (2022); Murphy et al. (2020); Mandal et al. (2021), which have been attributed to complex material memory effects Keim et al. (2019).
The notion of a distribution of local stresses in an amorphous material, rather than a single macroscopic average stress, has a well developed history in the statistical physics literature. Modelling approaches based on shear transformation zones Falk and Langer (1998) and lattice elastoplastic descriptions are based upon it Hébraud and Lequeux (1998); Nicolas et al. (2018); Bocquet et al. (2009); Lin et al. (2014), as is the SGR model itself Sollich et al. (1997); Sollich (1998). These approaches in turn draw on concepts originally developed by Spaepan Spaepen (1977) and Argon Argon and Kuo (1979) in the context of metallic glasses and bubbles rafts. More recently, frustrated local stresses have been recognised to play an important role in the elastically driven aging of soft solids Bouzid et al. (2017), the memory of soft jammed solids to shear flow Vinutha et al. (2024), the elasticity properties of amorphous solids Zhang et al. (2022), the directional memory of soft glasses Edera et al. (2025), and the non-monotonic relaxation of shear stress after flow cessation Ward and Fielding (2025). A key contribution of this work is to recognise and explore the significance of a stress distribution in strain recovery after unloading, and to emphasise that important effects such as non-monotonic recovery can be captured only in constitutive descriptions that indeed consider a full distribution of local stresses (or at least multiple moments of such a distribution), rather than a single average stress.
@@
The paper is structured as follows. In Sec. II we outline the SGR model Sollich et al. (1997); Sollich (1998), within which we shall study the phenomenon of strain recovery. The step stress protocol that we shall simulate is specified in Sec. III. In Sec. IV we define our units, summarise the parameters of the model and protocol, and discuss the rheological quantities to be measured. Our results are presented in Sec. V. We set out our conclusions in Sec. VIII.
@@ II Soft glassy rheology model
The soft glassy rheology (SGR) model Sollich et al. (1997); Sollich (1998) considers the dynamics of an ensemble of elastoplastic elements that explore a landscape of energy traps. Each element is notionally taken to represent a local mesoscopic region of a soft glassy material: a few tens of droplets in a dense emulsion, for example. Each element is assumed small relative to any macroscopic variations in the flow field, on the one hand, yet large enough to allow the definition of local continuum variables of shear strain $l$ and shear stress $kl$, on the other. Here $k$ is an elastic constant, assumed the same for all elements. The local strain variable $l$ is taken to describe the element’s state of elastic deformation relative to a locally undeformed equilibrium. Between local plastic yielding events, defined below, the strain of each element is assumed elastically to follow the macroscopically imposed strain rate $\dot{\gamma}$, such that $\dot{l}=\dot{\gamma}$.
Plasticity is incorporated by assuming that these local stresses are intermittently released by local plastic yielding events. In each such yielding event, a mesoscopic region of material is assumed suddenly to rearrange itself into a new local configuration, relative to a new state of local equilibrium. This is modelled by the corresponding representative elastoplastic element hopping between traps in the model’s energy landscape: at any instant in time, an element that lies in a trap of energy depth $E$ and has local shear strain $l$ is assumed to have a probability per unit time of hopping given by Sollich and Cates (2012)
$r(E,l)=\tau_{0}^{-1}{\rm min}\left\{1,\exp\left[(\tfrac{1}{2}kl^{2}-E)/x\right]\right\},$ (1)
@@
in which $x_{\rm g}$ is the model’s glass transition (effective) temperature, discussed further below. The element is further assumed to reset its local strain $l$ to zero, so relaxing its stress. (At the end of Sec. V, we shall return to consider a distribution of post-hop strain values with non-zero width $l_{p}$, to model local frustration.)
@@
The advection term on the left hand side captures the elastic loading of each element within its trap. The first term on the right hand side describes the local plastic yielding events described above. Because these are modelled by hops out of traps, this takes the form of a ‘death’ term. The second term on the right hand side is a ‘rebirth’ term, describing the hopping of freshly yielded elements into the bottom of new traps, $l=0$, with the new trap depth selected randomly from the prior $\rho(E)$ of Eqn. 2. The ensemble average hopping rate
@@ III Protocol
Sample preparation prior to shear is modelled by assuming a sudden deep quench at some time $t=-t_{\rm w}$, from infinite effective temperature to the final working value $x$, which is held constant thereafter. For times $-t_{\rm w}<t<0$ the sample is allowed to age undisturbed, without any strain or stress being applied. Throughout most of the paper, we assume that all elemental strains are equal to zero during this initial ageing stage, $l=0$. An exception can be found in Figs. 3 and 4, where we assume an initial Gaussian distribution of $l$ values, of narrow standard deviation $l_{0}=0.05$. This does not affect the basic physics of interest in this work, but merely allows us to show as a narrow Gaussian in Fig. 4 what would otherwise be an unplottable initial delta function distribution. We return at the end of Sec. V to consider in more detail the effect of non-zero $l_{0}$ values on the physics that we report.
@@ IV Units, parameter values and measured quantities
To solve the SGR model numerically, (at least) two different methods can be used. In the first, one derives integral equations from the governing master equation, Eqn. 3, then uses numerical mathematics to solve these Sollich (1998). In the second method, which we adopt here, one instead directly simulates the stochastic hopping dynamics of a population of $M$ SGR elements. In any timestep $dt$, each element has its strain shifted elastically by an amount $l\to l+\dot{\gamma}(t)dt$, where $\dot{\gamma}(t)$ is the strain rate at that time. Any element also plastically hops into a new trap with probability $rdt$, with $r$ that element’s probability per unit time of hopping as given by Eqn. 1 above. In the limit of many elements $M\to\infty$, one recovers the solution of the governing equations. During any intervals in which the stress is held constant, the strain rate is calculated as $\dot{\gamma}=\langle lr\rangle_{P}$. This ensures that the rate of stress loss via plastic hopping between traps is exactly counterbalanced by stress gain via elements shifting their local strains within their existing traps. At the instant of any step stress of size $\Delta\sigma$, each element is simply shifted elastically in its trap by an amount $l\to l+\Delta\sigma/k$.
@@
At the instant of stress switch-off, all elements instantaneously strain elastically backwards, giving a global strain change $\Delta\gamma=-\sigma_{0}$ (in our units). We then measure the subsequent strain evolution as a function of the time $t-t_{\rm stop}^{+}$ since switch-off. Our interest in particular will be in the total reverse strain $\Delta\gamma_{\rm rec}$ accumulated in the long time limit, $t-t_{\rm stop}\to\infty$, beyond that which arose in the fast elastic recoil during the reverse step stress. As noted above, in rare cases we shall find $\Delta\gamma_{\rm rec}<0$: the further straining that arises after the initial reverse elastic recoil in fact takes place in the forward direction.
@@ V.1 Forward strain during stress application
Following the application of a step stress of magnitude $\sigma_{0}$ at time $t=0$, a material’s creep response is characterised by the strain curve $\gamma(t)$. As noted above, in the SGR model this shows an initially elastic step strain of magnitude $\gamma_{0}=\sigma_{0}/k\;(=\sigma_{0}$ in our units) at $t=0$, followed by subsequent plastic straining for $t>0$. Fig. 1a) shows this plastic part of the creep, beyond the initial elastic step, normalised by the stress amplitude $\sigma_{0}$. Results are shown for a fixed sample age $t_{\rm w}$, and several imposed stresses in curves upwards, spanning values from below the SGR model’s dynamic yield stress $\sigma_{\rm Y}$ to above it. (At the noise temperature $x=0.3$ considered in this work, the yield stress $\sigma_{\rm Y}=0.758$.) The corresponding strain rate curves $\dot{\gamma}(t)$ are shown in panel b).
For imposed stresses $\sigma_{0}<\sigma_{\rm Y}$, the strain rate decreases perpetually as a power law $\dot{\gamma}\sim t^{-\alpha}$. The exponent $10<\alpha<$ 1 depends on the noise temperature $x$. The strain correspondingly increases sublinearly as $\gamma\sim t^{1-\alpha}$, corresponding to sustained power-law creep. In contrast, for imposed stresses $\sigma_{0}>\sigma_{\rm Y}$, an initial interval of creep (for $\sigma_{0}\gtrapprox\sigma_{\rm Y}$ at least) is interrupted by yielding, in which the strain rate curves upwards before finally settling to a time-independent value prescribed by the steady state flow curve $\sigma_{\rm ss}(\dot{\gamma})$, with $\sigma_{\rm ss}-\sigma_{\rm Y}(x)\sim\dot{\gamma}^{1-x}$. The time delay before yielding scales as the sample age $t_{\rm w}$, with a prefactor that diverges as $\sigma_{0}-\sigma_{\rm Y}\to 0^{+}$ Fielding et al. (2000).
@@
Fig. 2b) shows the same data as in Fig. 1, but with the scaled strain rate $\dot{\gamma}(t)/\sigma_{0}$ now plotted parametrically as a function of the forward plastic strain $(\gamma(t)-\gamma_{0})/\sigma_{0}$. Accordingly, time $t$ is not explicitly represented in Fig. 2. The scaled strain on the abscissa however increases monotonically with increasing $t$. The counterpart data for sample ages $t_{\rm w}=10^{1}$ and $10^{6}$ are shown in Figs. 2a) and c) respectively. As can be seen, for imposed stresses above the yield stress, $\sigma_{0}>\sigma_{\rm Y}$, yielding typically occurs at a scaled strain $(\gamma-\gamma_{0})/\sigma_{0}=O(1)$, although the value increases slowly with $t_{\rm w}$ and diverges as $\sigma_{0}\to\sigma_{\rm Y}^{+}$.
A dot at each of several logarithmically spaced values of the abscissa in Fig. 2 shows the values of scaled forward plastic strain $(\gamma-\gamma_{0})/\sigma_{0}=\Delta\gamma_{\rm f}/\sigma_{0}$ at which we shall later, in Fig. 5, explore the amount of strain $\Delta\gamma_{\rm rec}$ that is ultimately recovered plastically after the stress is switched off.
@@ V.2 Strain recovery after stress switch-off
Having mapped out the forward creep and/or yielding and flow that arise while the stress is held imposed, we now explore the strain recovery after the stress is switched off. Our focus will be in particular on the part of this recovery that stems from slow ongoing plastic yielding of elements that were left with negative local stresses as a result of the fast initial elastic recoil. In Sec. VI below, we shall consider the interplay of a non-zero timescale for the initial (visco)elastic recoil with the later slow plastic strain recovery.
@@
The basic physics that we aim to elucidate is shown in Figs. 3 and 4. The imposed stress $\sigma$ as a function of time $t$ is shown in the top panel of Fig. 3. The corresponding strain response $\gamma(t)$ is shown in the bottom panel. As described above, a stress of amplitude $\sigma_{0}$ is switched on at time $t=0$. (The preparation phase during which the sample is allowed to age undisturbed for a time $t_{\rm w}$ prior to $t=0$ is not shown.) At the instant of switch-on, the system shows an instantaneous forward elastic step strain of amplitude $\gamma_{0}=\sigma_{0}$, in our units in which $k=1$. (In experimental practice, this forward strain would be rapid but not instantaneous, with a rate limited by inertia or dissipation; we return to this point in Sec. VI below.) The stress is then held constant for some time interval, over which the strain increases further due to plastic yielding events, this forward straining having been explored in Sec. V.1. Once it has thus increased by an amount $\Delta\gamma_{\rm f}$ beyond $\gamma_{0}$, the stress is switched off. This defines the switch-off time $t_{\rm stop}$. At the instant of switch-off, $t=t_{\rm stop}$, the system shows an instantaneous backward elastic step strain, $\Delta\gamma=-\sigma_{0}=-\gamma_{0}$. (We consider in Sec. VI the effect of a finite timescale for this initial (visco)elastic recoil.) The stress is then held at zero for all subsequent times $t>t_{\rm stop}$. As a function of the time interval $t-t_{\rm stop}^{+}$, the strain further recovers in the negative direction by an amplitude $\Delta\gamma_{\rm rec}$, beyond the negative elastic step $\Delta\gamma=-\gamma_{0}$ that occurred at the instant of switch-off. This additional recovery $\Delta\gamma_{\rm rec}$, which has a significant amplitude compared with both $\Delta\gamma_{\rm f}$ and $\gamma_{0}$, arises from slowly ongoing plastic events, in which elements left with negative local stresses by the initial elastic recoil slowly reset these negative stresses to zero.
To understand the physics underlying this strain evolution in more detail, let us consider the master equation governing the time evolution of the joint probability distribution of elemental trap depths $E$ and strains $l$. This was specified in Eqn. 3, which we repeat here for convenience:
@@
In contrast, during any interval in which the stress is constant, $\dot{\sigma}=0$, either after switch-on while the stress of amplitude $\sigma=\sigma_{0}$ is held imposed, or after switch-off when $\sigma=0$, the strain rate is given by
@@
This tells us that any forward straining $\Delta\gamma_{\rm f}$ that occurs after the initial elastic loading during switch-on but before switch-off is determined purely by the relaxation of local stresses due to local plastic yielding events. Importantly, it also tells us that any strain recovery $\Delta\gamma_{\rm rec}$ that occurs after the initial elastic recoil post switch-off is likewise determined by the slow plastic yielding of elements left with negative local stress values by that initial recoil. Indeed, as these local plastic events occur and elements hop between traps, a global straining must arise in order to advect elements within their traps and exactly counterbalance any such plastic stress loss (or gain), keeping the overall stress constant. Accordingly, the initial fast elastic recoil creates a state out of which the slow reverse plastic events later emerge, and we return in Sec. VI below to consider in more detail the interplay between these two different stages of strain recovery.
In order fully to understand the forward plastic straining during stress imposition and the strain recovery afterwards, therefore, we must consider in more detail the evolution of the probability distribution of local strains, and the way this determines the strain rate $\dot{\gamma}(t)$ via Eqn. 9. Accordingly, we plot this distribution $P(l,t)$ in Fig. 4 (having first integrated out the trap energy depth variable $E$). We do so at several times $t$ while the stress is held applied (top panel in Fig. 4), and at several times after the stress is switched off (bottom panel). The time at which any distribution is plotted in Fig. 4 is indicated by a cross of corresponding colour in Fig. 3.
@@
Were the strain held constant, with zero strain rate $\dot{\gamma}=0$, the macroscopic consequence of these plastic yielding events would be a relaxation of the ensemble average stress at rate $\dot{\sigma}=\langle l\,r(E,l)\rangle_{P(t)}$. Conversely, to maintain the stress constant, $\sigma=+\sigma_{0}$ with $\dot{\sigma}=0$, requires a forward strain of corresponding rate $\dot{\gamma}=\langle l\,r(E,l)\rangle_{P(t)}$, ensuring sufficient forward advection of elements within their traps exactly to counterbalance this plastic relaxation. This is the origin of the forward plastic straining seen for times $0<t<t_{\rm stop}$ in Fig. 3. Its consequence is to advect forward at rate $\dot{l}=\dot{\gamma}$ both the spike in probability centred (originally) at $l=\sigma_{0}=\gamma_{0}$ (as it simultaneously depletes due to the plastic relaxation events just described), as well as the newly developing lobe of probability near the origin $l\approx 0$. In this way, the distribution evolves over time from the black to cyan curve in the top panel of Fig. 4.
At the instant the stress is switched off, a fast elastic recoil occurs, in which the distribution of local strains instantaneously shifts by $l\to l-\gamma_{0}$, with $\gamma_{0}=\sigma_{0}$, without changing shape, as each element elastically reduces its strain within its trap. Accordingly, the rightmost (cyan) curve in the top panel of Fig. 4, which obtained at $t=t_{\rm stop}^{-}$, shifts to become the rightmost (magenta) curve in the bottom panel at $t=t_{\rm stop}^{+}$.